# Multi-view Disentanglement for Reinforcement Learning with Multiple Cameras

### v1

@@ Abstract
The performance of image-based Reinforcement Learning (RL) agents can vary depending on the position of the camera used to capture the images. Training on multiple cameras simultaneously, including a first-person egocentric camera, can leverage information from different camera perspectives to improve the performance of RL. However, hardware constraints may limit the availability of multiple cameras in real-world deployment. Additionally, cameras may become damaged in the real-world preventing access to all cameras that were used during training. To overcome these hardware constraints, we propose Multi-View Disentanglement (MVD), which uses multiple cameras to learn a policy that achieves zero-shot generalisation to any single camera from the training set. Our approach is a self-supervised auxiliary task for RL that learns a disentangled representation from multiple cameras, with a shared representation that is aligned across all cameras to allow generalisation to a single camera, and a private representation that is camera-specific. We show experimentally that an RL agent trained on a single third-person camera is unable to learn an optimal policy in many control tasks; but, our approach, benefiting from multiple cameras during training, is able to solve the task using only the same single third-person camera.
@@ 1 Introduction
The ability of a Reinforcement Learning (RL) agent to learn an optimal policy on robotic control tasks from images depends on the position of the camera available during training. Often, a static third-person camera pointing towards the scene (e.g. Figure 1(b)) is not sufficient to learn an optimal policy. A first-person egocentric camera on the robot’s end-effector (e.g. Figure 1(a)) has been shown to be necessary to learn an optimal policy in many tasks (Hsu et al., 2022). Approaches that leverage both first-person and third-person cameras simultaneously have been shown to improve the performance of RL algorithms (Hsu et al., 2022; Jangir et al., 2022; Barati & Chen, 2019). Whilst it is possible to create multiple camera views in simulation to improve training, access to multiple cameras in the real-world may be restricted due to hardware limitations. Therefore, it is desirable for an RL agent to be able to generalise from multiple cameras to a single camera for successful deployment. Even when multiple cameras are available in the real-world, cameras may fail or become faulty during deployment, so a robust RL agent should continue to perform optimally with only a subset of the cameras that were available during training. We propose to address these hardware limitations with an approach that leverages multiple cameras during training to learn a policy that successfully performs control tasks with multiple cameras as well as achieving zero-shot generalisation to a single camera.
@@
We use multiple cameras to learn a disentangled representation that allows zero-shot generalisation to any one of the training cameras. Our approach, called Multi-View Disentanglement (MVD), is a self-supervised auxiliary task to learn such a disentangled representation and can be used with existing RL algorithms. The learned representation is disentangled into a shared representation that is similar across all cameras, and a private representation that contains information only available to an individual camera. The shared representation gives the RL policy a consistent representation that can be relied upon regardless of the camera. The private representation allows information only available to a single camera to be used during training to improve policy learning as some camera views make it easier to discern important features than others.
@@
We evaluate our approach on robotic control tasks using a Panda robot (Gallouédec et al., 2021; Hsu et al., 2022), and several MetaWorld tasks (Yu et al., 2020) using a Sawyer robot. We show experimentally that an RL agent often cannot learn an optimal policy when training on a third-person camera alone, and that an approach combining multiple cameras during training (Hsu et al., 2022) does not generalise to a single camera. Our results also show that our approach, MVD, is able to learn an optimal policy from multiple cameras and achieve zero-shot generalisation to successfully solve the task using any of the cameras individually in many tasks. We also analyse the learned representation of our approach using saliency maps and show that the shared representation focuses on the key features in the image visible in all cameras, such as the position of the goal, whereas the private representation focuses on features only visible in a single camera.
@@ 2 Related work
Prior work uses multiple cameras to improve performance on robotic control tasks. Hsu et al. (2022) use both first-person and third-person cameras together with a variational information bottleneck to regularise the third-person representation, Jangir et al. (2022) use transformers with cross-view attention, and Barati & Chen (2019) train multiple workers with different views and combine features for each worker weighted by Q-values. These approaches combine camera representations to learn a policy that is dependent on all available cameras, and so cannot generalise if one of the training cameras is no longer available. Acar et al. (2023) trains a teacher RL policy on multiple cameras augmented with human demonstrations. They use imitation learning to learn a student single-camera policy to output an action similar to the teacher multi-camera policy. Our approach does not require training separate teacher and student polices, and instead learns a disentangled representation online that allows generalisation to a single camera. We also do not require human demonstrations during training.
@@
Disentangled representations have been used to improve RL generalisation for a single camera (Dunion et al., 2023a; b; Higgins et al., 2017). Other approaches learn representations using mutual information (Lee et al., 2020), regression targets (McInroe et al., 2024) and similarity constraints (Agarwal et al., 2021; Mazoure et al., 2020; van den Oord et al., 2018). CURL (Laskin et al., 2020) uses contrastive learning to maximise the similarity between representations of the same image with different augmentations to improve sample efficiency. We also use contrastive learning with images, but consider camera views rather than augmentations.
@@ 4 Multi-view disentanglement for RL
We propose Multi-View Disentanglement (MVD) to learn a disentangled representation from multiple cameras that allows zero-shot generalisation to a single camera. Our approach learns separate shared $\mathbf{s}_{t}^{c_{i}}$ and private $\mathbf{p}_{t}^{c_{i}}$ representations for each camera $c_{i}\in C$, where $C$ denotes the set of cameras and $t$ is the timestep. The shared representation $\mathbf{s}_{t}^{c_{i}}$ is trained to be similar across all cameras $\forall c_{i}\in C$, while the private representation $\mathbf{p}_{t}^{c_{i}}$ encodes information available only to a specific camera $c_{i}\in C$. Both representations together are used to condition the RL policy $\pi(\mathbf{a}_{t}|\mathbf{z}_{t})$ with $\mathbf{z}_{t}=(\mathbf{s}_{t}^{c_{i}},\mathbf{p}_{t}^{c_{j}})$, where the shared part allows for camera generalisation and the private part allows the policy to leverage extra information available to a single camera to improve training. We define two contrastive learning objectives that together give an auxiliary task for multiple camera views that can be applied to existing RL algorithms. We first provide an overview of the approach in Section 4.1, then discuss the MVD auxiliary task in Section 4.2. Finally, we explain how the disentangled representation is used as input to the RL policy in Section 4.3.
@@ 5.1 Experimental setup
To demonstrate the broad applicability of our approach, we apply MVD to two RL algorithms and different numbers of cameras. For the Panda tasks we use SAC (Haarnoja et al., 2018) with a decoder to aid learning from images and three cameras (depicted in Figure 3). For the MetaWorld tasks we use DrQ (Yarats et al., 2021a), which uses image augmentations to improve sample efficiency on these more difficult tasks, and two cameras (depicted in Figure 1). In all tasks, one of the cameras is a first-person camera on the end-effector, while the remaining views are static third-person cameras. Implementation details and hyperparameters are provided in Appendix B.1.
@@ 5.2 Generalisation results
The ablation experiment of MVD with the shared representation only (MVD-SharedOnly) demonstrates the importance of the private representation in MVD. In all the tasks except Panda Reach, the shared representation alone achieves lower performance than MVD. This may be because the first-person camera is easier to learn from than the third-person cameras (as evidenced by the single camera baseline). Encouraging the shared representations to be similar for all cameras, without the flexibility of a separate private representation, may prevent important features from a single camera being used during training if the agent is not yet able to extract similar features from the other cameras.
5.3 Representation analysis
Panda Reach MetaWorld Soccer
                              (front)             (side)
Images
Refer to caption Refer to caption
Shared
Refer to caption Refer to caption
Private
Refer to caption Refer to caption
We use saliency maps to show that our learned representations reflect shared and private features in the camera images. The attribution method we use is Integrated Gradients (Sundararajan et al., 2017) with SmoothGrad-Squared (Hooker et al., 2018) to reduce visual noise. Using the learned encoders $f_{\theta}$ and $g_{\phi}$, we calculate the attributions for each image pixel on each representation feature. We also calculate the attributions for each dimension in the representations on the output of the learned policy $\pi$. The pixel attributions are weighted by the corresponding policy attributions, and normalised for each camera to be in $[0,1]$, to visualise the attention of the RL policy based on the features in the shared and private representations separately. Further details of the attribution calculations are provided in Appendix B.2.
@@ 6 Conclusion and future work
We demonstrated that camera perspective impacts the ability of an RL agent to learn an optimal policy. The impact can be mitigated by training with multiple cameras in simulation, but hardware constraints may prevent an RL agent from always relying on access to all of these cameras in the real-world. We propose Multi-View Disentanglement (MVD), an auxiliary task for RL algorithms to learn disentangled representations with a shared representation, which is aligned across all cameras, and a private representation, which is camera-specific. Our experiments showed that an RL agent trained only on a single third-person camera cannot learn an optimal policy in many control tasks, whereas MVD, benefiting from multiple cameras in training, achieves zero-shot generalisation to solve the task with same single third-person camera.
@@ B.1 MVD implementation
B.2 Saliency maps
@@
Appendix C Environment details

### v2

@@ Abstract
The performance of image-based Reinforcement Learning (RL) agents can vary depending on the position of the camera used to capture the images. Training on multiple cameras simultaneously, including a first-person egocentric camera, can leverage information from different camera perspectives to improve the performance of RL. However, hardware constraints may limit the availability of multiple cameras in real-world deployment. Additionally, cameras may become damaged in the real-world preventing access to all cameras that were used during training. To overcome these hardware constraints, we propose Multi-View Disentanglement (MVD), which uses multiple cameras to learn a policy that is robust to a reduction in the number of cameras to generalise to any single camera from the training set. Our approach is a self-supervised auxiliary task for RL that learns a disentangled representation from multiple cameras, with a shared representation that is aligned across all cameras to allow generalisation to a single camera, and a private representation that is camera-specific. We show experimentally that an RL agent trained on a single third-person camera is unable to learn an optimal policy in many control tasks; but, our approach, benefiting from multiple cameras during training, is able to solve the task using only the same single third-person camera.
@@ 1 Introduction
The ability of a Reinforcement Learning (RL) agent to learn an optimal policy on robotic control tasks from images depends on the position of the camera available during training. Often, a static third-person camera pointing towards the scene (e.g. Figure 1(b)) is not sufficient to learn an optimal policy. A first-person egocentric camera on the robot’s end-effector (e.g. Figure 1(a)) has been shown to be necessary to learn an optimal policy in many tasks (Hsu et al., 2022). Approaches that leverage both first-person and third-person cameras simultaneously have been shown to improve the performance of RL algorithms (Hsu et al., 2022; Jangir et al., 2022; Barati & Chen, 2019). Whilst it is possible to create multiple camera views in simulation to improve training, access to multiple cameras in the real-world may be restricted due to hardware limitations. Therefore, it is desirable for an RL agent to be able to generalise from multiple cameras to a single camera for successful deployment. Even when multiple cameras are available in the real-world, cameras may fail or become faulty during deployment, so a robust RL agent should continue to perform optimally with only a subset of the cameras that were available during training. We propose to address these hardware limitations with an approach that leverages multiple cameras during training to learn a policy that successfully performs control tasks with multiple cameras as well as achieving robustness to a reduction in the available cameras.
@@
We use multiple cameras to learn a disentangled representation that allows robustness to a reduction in cameras to generalise to any single camera from the training set. Our approach, called Multi-View Disentanglement (MVD), is a self-supervised auxiliary task to learn such a disentangled representation and can be used with existing RL algorithms. The learned representation is disentangled into a shared representation that is similar across all cameras, and a private representation that contains information only available to an individual camera. The shared representation gives the RL policy a consistent representation that can be relied upon regardless of the camera. The private representation allows information only available to a single camera to be used during training to improve policy learning as some camera views make it easier to discern important features than others.
@@
We evaluate our approach on robotic control tasks using a Panda robot (Gallouédec et al., 2021; Hsu et al., 2022), and several MetaWorld tasks (Yu et al., 2020) using a Sawyer robot. We show experimentally that an RL agent often cannot learn an optimal policy when training on a third-person camera alone, and that an approach combining multiple cameras during training (Hsu et al., 2022) does not generalise to a single camera. Our results also show that our approach, MVD, is able to learn an optimal policy from multiple cameras and achieve zero-shot generalisation to successfully solve the task using any of the cameras individually in many tasks.
@@ 2 Related work
Prior work uses multiple cameras to improve performance on robotic control tasks. Hsu et al. (2022) use both first-person and third-person cameras together with a variational information bottleneck to regularise the third-person representation, Jangir et al. (2022) use transformers with cross-view attention, Barati & Chen (2019) train multiple workers with different views and combine features for each worker weighted by Q-values, and Driess et al. (2022) use Neural Radiance Fields to learn a representation from multiple images to improve the performance of RL. These approaches combine camera representations to learn a policy that is dependent on all available cameras, and so cannot generalise if one of the training cameras is no longer available. Acar et al. (2023) trains a teacher RL policy on multiple cameras augmented with human demonstrations. They use imitation learning to learn a student single-camera policy to output an action similar to the teacher multi-camera policy. Shang & Ryoo (2021) also use imitation learning where demonstrations are a third-person view from a human or robot, which are used alongside first-person views of the egocentric robot to learn a disentangled representation for imitation learning. Our approach does not require training a separate teacher policy or collecting suitable demonstrations, and instead learns a disentangled representation online that allows generalisation to a single camera.
@@
Disentangled representations have been used to improve RL generalisation for a single camera (Dunion et al., 2023a; b; Higgins et al., 2017). Other approaches learn representations using mutual information (Lee et al., 2020; Garcin et al., 2024), regression targets (McInroe et al., 2024) and similarity constraints (Agarwal et al., 2021; Mazoure et al., 2020; van den Oord et al., 2018). CURL (Laskin et al., 2020) uses contrastive learning to maximise the similarity between representations of the same image with different augmentations to improve sample efficiency. However, Li et al. (2022) find that self-supervised learning frameworks with augmented images have limited impact on RL performance compared to image augmentation alone. While we also use contrastive learning with images, we consider camera views rather than augmentations.
@@ 4 Multi-view disentanglement for RL
We propose Multi-View Disentanglement (MVD) to learn a disentangled representation from multiple cameras that is robust to a reduction in available cameras. Our approach learns separate shared $\mathbf{s}_{t}^{c_{i}}$ and private $\mathbf{p}_{t}^{c_{i}}$ representations for each camera $c_{i}\in C$, where $C$ denotes the set of cameras and $t$ is the timestep. The shared representation $\mathbf{s}_{t}^{c_{i}}$ is trained to be similar across all cameras $\forall c_{i}\in C$, while the private representation $\mathbf{p}_{t}^{c_{i}}$ encodes information available only to a specific camera $c_{i}\in C$. Both representations together are used to condition the RL policy $\pi(\mathbf{a}_{t}|\mathbf{z}_{t})$ with $\mathbf{z}_{t}=(\mathbf{s}_{t}^{c_{i}},\mathbf{p}_{t}^{c_{j}})$, where the shared part allows for camera generalisation and the private part allows the policy to leverage extra information available to a single camera to improve training. We define two contrastive learning objectives that together give an auxiliary task for multiple camera views that can be applied to existing RL algorithms. We first provide an overview of the approach in Section 4.1, then discuss the MVD auxiliary task in Section 4.2. Finally, we explain how the disentangled representation is used as input to the RL policy in Section 4.3.
@@ 5.1 Experimental setup
To demonstrate the broad applicability of our approach, we apply MVD to two RL algorithms and different numbers of cameras. For the Panda tasks we use SAC (Haarnoja et al., 2018) with a decoder to aid learning from images and three cameras (depicted in Figure 3). For the MetaWorld tasks we use DrQ (Yarats et al., 2021a), which uses image augmentations to improve sample efficiency on these more difficult tasks, and two cameras (depicted in Figure 1). In all tasks, one of the cameras is a first-person camera on the end-effector, while the remaining views are static third-person cameras. Implementation details and hyperparameters are provided in Appendix B.1. Our code is available at github.com/uoe-agents/MVD.
@@ 5.2 Generalisation results
The ablation of MVD with the shared representation only (MVD-SharedOnly) demonstrates the importance of the private representation in MVD. In all the tasks except Panda Reach, the shared representation alone achieves lower performance than MVD. This may be because the first-person camera is easier to learn from than the third-person cameras (as evidenced by the single camera baseline). Encouraging the shared representations to be similar for all cameras, without the flexibility of a separate private representation, may prevent important features from a single camera being used during training if the agent is not yet able to extract similar features from the other cameras.
5.3 Ablation study
In this section, we conduct more detailed analysis of MVD for the Panda Reach and MetaWorld Soccer tasks with ablation experiments to better understand the components of MVD. We also provide some analysis of the learned representations in Appendix C.
all cameras first-person third-person (front) third-person (side)
Reach
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Results of ablation experiments for Panda Reach showing success rate for evaluation on all cameras (left of dashed line) compared with success rate on each of the individual cameras (right). Success rate is averaged over 20 evaluation episodes for 5 seeds. The shaded region is standard deviation. See text for details of each ablation.
all cameras first-person third-person
Soccer
Refer to caption Refer to caption Refer to caption
Figure 7: Results of ablation experiments for MetaWorld Soccer showing success rate for evaluation on all cameras (left of dashed line) compared with success rate on each of the individual cameras (right). Success rate is averaged over 20 evaluation episodes for 5 seeds. The shaded region is standard deviation. See text for details of each ablation.
We consider the MVD shared loss (Equation 3), which uses two types of negative samples: shared representations from the same camera at different timesteps, and private representations for all cameras at the same timestep. We assess the impact of each type of negative sample by measuring performance when one type is removed, resulting in two MVD ablations: MVD shared loss without shared negatives and MVD shared loss without private negatives. The results in Figures 6 and 7 show that removing either of the negative samples reduces the performance of MVD, but the shared negatives have a much greater impact on performance than the private negatives. This may be due to the other constraints in the MVD shared and private losses indirectly encouraging disentanglement by encouraging shared representations to be similar and private representations to be dissimilar, limiting the additional improvement gained by explicitly doing so through the use of private negatives in the shared loss.
To understand the impact of the similarity measure for InfoNCE, we compare the normalised dot product similarity measure used for MVD to an alternative bilinear similarity measure. The results, also in Figures 6 and 7, show that while the method is reasonably robust to the choice of similarity measure, the dot product outperforms the bilinear similarity measure, particularly in Panda Reach. This may be due to the learnable weight matrix in the bilinear similarity measure making it a more lenient constraint as the agent can learn to give less weight to some dimensions than others.
Finally, we consider the same training setup as MVD but without the MVD loss. We train the base algorithm with randomised cameras to replicate the MVD training setup by randomly sampling a camera representation as input to the RL loss at each timestep (as described in Section 4.3) but without the MVD loss to structure the learned representation. As in the previous section, the base algorithm is SAC for Panda Reach and DrQ for MetaWorld Soccer. The results in Figures 6 and 7 show that randomising the cameras alone is not enough to learn to solve the more difficult MetaWorld Soccer task. While camera randomisation does improve on the single camera baseline (see Figure 4) for the easier Panda Reach task, it achieves a much lower success rate than MVD.
@@ 6 Conclusion and future work
We demonstrated that camera perspective impacts the ability of an RL agent to learn an optimal policy. The impact can be mitigated by training with multiple cameras in simulation, but hardware constraints may prevent an RL agent from always relying on access to all of these cameras in the real-world. We propose Multi-View Disentanglement (MVD), an auxiliary task for RL algorithms to learn disentangled representations with a shared representation, which is aligned across all cameras, and a private representation, which is camera-specific. Our experiments showed that an RL agent trained only on a single third-person camera cannot learn an optimal policy in many control tasks, whereas MVD, benefiting from multiple cameras in training, achieves robustness to a reduction in cameras to solve the task with the same single third-person camera.
@@ B.1 MVD implementation
Appendix C Representation analysis
We use saliency maps to show that our learned representations reflect shared and private features in the camera images. The attribution method we use is Integrated Gradients (Sundararajan et al., 2017) with SmoothGrad-Squared (Hooker et al., 2018) to reduce visual noise. However, while features attribution methods such as Integrated Gradients are commonly used in the literature, very recently Bilodeau et al. (2024) has shown that such feature attributions methods can fail to improve on random guessing for inferring model behaviour. We provide our representation analysis results for completeness.
C.1 Saliency map results
Panda Reach MetaWorld Soccer
                              (front)             (side)
Images
Refer to caption Refer to caption
Shared
Refer to caption Refer to caption
Private
Refer to caption Refer to caption
Using the learned encoders $f_{\theta}$ and $g_{\phi}$, we calculate the attributions for each image pixel on each representation feature. We also calculate the attributions for each dimension in the representations on the output of the learned policy $\pi$. The pixel attributions are weighted by the corresponding policy attributions, and normalised for each camera to be in $[0,1]$, to visualise the attention of the RL policy based on the features in the shared and private representations separately.
C.2 Implementation details
@@
Appendix D Environment details